OpenStudy (yrivers36):

solve the rational equation: x/x-2=2x-2/x-2

OpenStudy (anonymous):

Ok Since they have a common denominator of (x-2), if you were to multiply both sides by (x-2)/1 you would cancel the denominator away. This leaves you with x = 2x-2 0 = x-2 x = 2

OpenStudy (yrivers36):

i did get 2 but was not sure..thanx again

OpenStudy (anonymous):

You're welcome